# A quantum implementation of high-order power method for estimating geometric entanglement of pure states

### v1

@@ 3.3 Mitigating error for arbitrary noise models
To mitigate the results of our algorithm on hardware or simulations where the noise model is usually unknown, we use the following procedure. We run the algorithm on a reference state with a known $\widetilde{E}_{G}$ and measure the noisy value $E_{G}$. In our case we chose GHZ[$n$] with $\widetilde{E}_{G}=0.50$ since the value is consistent for any number of qubits. Using this value and Eq. (25) we calculate $p$ as follows:
$p=1-\left(\frac{1-\widetilde{E}_{G}}{1-\overline{E}_{G}}\right)^{2d}.$ (26)
@@ A.1 Implementation of Gauss-Seidel method and Shifted HOPM algorithms
${\mathbf{v}}_{i}^{(k+1)}=\lambda^{(k)}\psi\times_{\hat{i}}\left({\mathbf{v}_{1}^{(k+1)}}^{\ast},\ldots,{\mathbf{v}_{i-1}^{(k+1)}}^{\ast},{\mathbf{v}_{i+1}^{(k)}}^{\ast},\ldots,{\mathbf{v}_{n}^{(k)}}^{\ast}\right)+\beta{\mathbf{v}}_{i}^{(k)},$ (27)
@@
To implement GSM the only change to our algorithm is in the classical step where we recover the angles $(\tilde{\vartheta}_{i}^{(k+1)},\tilde{\varphi}_{i}^{(k+1)})$ using one-qubit tomography. Instead of Eq. (9) we need to solve the following system of equations to find $({\vartheta}_{i}^{(k+1)},{\varphi}_{i}^{(k+1)})$,
$A^{2}$ $\cos^{2}\frac{\vartheta_{i}^{(k+1)}}{2}=\left|\lambda^{(k)}e^{-i\tilde{\varphi}_{i}^{(k+1)}/2}\cos\frac{\tilde{\vartheta}_{i}^{(k+1)}}{2}\right.+\left.\beta e^{-i\varphi_{i}^{(k)}/2}\cos\frac{\vartheta_{i}^{(k)}}{2}\right|^{2};$ (28)
$A^{2}$ $\sin\vartheta_{i}^{(k+1)}\sin\varphi_{i}^{(k+1)}=(\lambda^{(k)})^{2}\sin\tilde{\vartheta}_{i}^{(k+1)}\sin\tilde{\varphi}_{i}^{(k+1)}+\beta^{2}\sin\vartheta_{i}^{(k)}\sin\varphi_{i}^{(k)}+2\beta\lambda^{(k)}\sin\frac{\tilde{\vartheta}_{i}^{(k+1)}+\vartheta_{i}^{(k)}}{2}\sin\frac{\tilde{\varphi}_{i}^{(k+1)}+\varphi_{i}^{(k)}}{2};$
$A^{2}$ $\sin\vartheta_{i}^{(k+1)}\cos\varphi_{i}^{(k+1)}=(\lambda^{(k)})^{2}\sin\tilde{\vartheta}_{i}^{(k+1)}\cos\tilde{\varphi}_{i}^{(k+1)}+\beta^{2}\sin\vartheta_{i}^{(k)}\cos\varphi_{i}^{(k)}+2\beta\lambda^{(k)}\sin\frac{\tilde{\vartheta}_{i}^{(k+1)}+\vartheta_{i}^{(k)}}{2}\cos\frac{\tilde{\varphi}_{i}^{(k+1)}+\varphi_{i}^{(k)}}{2},$
@@
$A^{2}=(\lambda^{(k)})^{2}+\beta^{2}+2\beta\lambda^{(k)}\cos\frac{\tilde{\vartheta}_{i}^{(k+1)}-\vartheta_{i}^{(k)}}{2}\cos\frac{\tilde{\varphi}_{i}^{(k+1)}-\varphi^{(k)}}{2}$
@@ A.2 Supplemental simulation results
Figure 7: Simulation and mitigation results for the QHOPM algorithm on the W[$3$] state. (a) shows the effect of noise on convergence. (b) shows the effects of mitigation on the noisy simulation. $\widetilde{E}_{G}=5/9\approx 0.554$. Each simulation was run 10 times (with the same 10 random initial separable states) with $1\times 10^{5}$ shots per each measurement. Solid lines represent the mean geometric entanglement, $E_{G}{}$, and the lightly shaded colours represent the standard deviation. The colour of the lines represents the noise model used for the simulation (see legend). Red dotted lines show results of classical HOPM.

### v2

@@ 3.3 Mitigating error for arbitrary noise models
To mitigate the results of our algorithm on hardware or simulations where the noise model is usually unknown, we use the following procedure. We run the algorithm on a reference state with a known $\hat{E}_{G}$ and measure the noisy value $E_{G}$. In our case we chose GHZ[$n$] with $\hat{E}_{G}=0.50$ since the value is consistent for any number of qubits. Using this value and Eq.  (26) we calculate $p$ as follows:
$p=1-\left(\frac{1-\hat{E}_{G}}{1-\overline{E}_{G}}\right)^{2d}.$ (27)
@@ A.1 Implementation of Gauss-Seidel method and Shifted HOPM algorithms
${\mathbf{v}}_{i}^{(k+1)}=\lambda^{(k)}\psi\times_{{i}}\left({\mathbf{v}_{1}^{(k+1)}}^{\ast},\ldots,{\mathbf{v}_{i-1}^{(k+1)}}^{\ast},{\mathbf{v}_{i+1}^{(k)}}^{\ast},\ldots,{\mathbf{v}_{n}^{(k)}}^{\ast}\right)+\beta{\mathbf{v}}_{i}^{(k)},$ (28)
@@
To implement GSM the only change to our algorithm is in the classical step where we recover the angles $(\check{\vartheta}_{i}^{(k+1)},\check{\varphi}_{i}^{(k+1)})$ using one-qubit tomography. Instead of Eq.  (10) we need to solve the following system of equations to find $({\vartheta}_{i}^{(k+1)},{\varphi}_{i}^{(k+1)})$,
$A^{2}\cos$ ${}^{2}\frac{\vartheta_{i}^{(k+1)}}{2}=\left|\lambda^{(k)}e^{-i\check{\varphi}_{i}^{(k+1)}/2}\cos\frac{\check{\vartheta}_{i}^{(k+1)}}{2}\right.+\left.\beta e^{-i\varphi_{i}^{(k)}/2}\cos\frac{\vartheta_{i}^{(k)}}{2}\right|^{2};$ (29)
$A^{2}\sin$ $\vartheta_{i}^{(k+1)}\sin\varphi_{i}^{(k+1)}=(\lambda^{(k)})^{2}\sin\check{\vartheta}_{i}^{(k+1)}\sin\check{\varphi}_{i}^{(k+1)}$
$+\beta^{2}\sin\vartheta_{i}^{(k)}\sin\varphi_{i}^{(k)}+2\beta\lambda^{(k)}\sin\frac{\check{\vartheta}_{i}^{(k+1)}+\vartheta_{i}^{(k)}}{2}\sin\frac{\check{\varphi}_{i}^{(k+1)}+\varphi_{i}^{(k)}}{2};$
$A^{2}\sin$ $\vartheta_{i}^{(k+1)}\cos\varphi_{i}^{(k+1)}=(\lambda^{(k)})^{2}\sin\check{\vartheta}_{i}^{(k+1)}\cos\check{\varphi}_{i}^{(k+1)}$
$+\beta^{2}\sin\vartheta_{i}^{(k)}\cos\varphi_{i}^{(k)}+2\beta\lambda^{(k)}\sin\frac{\check{\vartheta}_{i}^{(k+1)}+\vartheta_{i}^{(k)}}{2}\cos\frac{\check{\varphi}_{i}^{(k+1)}+\varphi_{i}^{(k)}}{2},$
@@
$A^{2}=(\lambda^{(k)})^{2}+\beta^{2}+2\beta\lambda^{(k)}\cos\frac{\check{\vartheta}_{i}^{(k+1)}-\vartheta_{i}^{(k)}}{2}\cos\frac{\check{\varphi}_{i}^{(k+1)}-\varphi^{(k)}}{2}$
@@ A.2 Supplemental simulation results
Figure 8: Simulation and mitigation results for the QHOPM algorithm on the W[$3$] state. (a) shows the effect of noise on convergence. (b) shows the effects of mitigation on the noisy simulation. $\hat{E}_{G}=5/9\approx 0.554$. Each simulation was run 10 times (with the same 10 random initial separable states) with $1\times 10^{5}$ shots per each measurement. Solid lines represent the mean geometric entanglement, $E_{G}{}$, and the lightly shaded colours represent the standard deviation. The colour of the lines represents the noise model used for the simulation (see legend). Red dotted lines show results of classical HOPM.